OpenStudy (anonymous):

Which of these is a simplified form of the equation 7x + 4 = 5 + 3x + 1x? (4 points) 11x = 1 3x = 9 11x = 9 3x = 1

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

7x + 4 = 5 + 3x + 1x First, combine 3x + 1x

OpenStudy (igreen):

3x + 1x is the same as 3 + 1 but we add an 'x' at the end.

OpenStudy (anonymous):

4x?

OpenStudy (igreen):

Yep! So we have: 7x + 4 = 5 + 4x Now subtract 4x to both sides, what's 7x - 4x?

OpenStudy (anonymous):

3x?

OpenStudy (anonymous):

oh so its 3x=9?

OpenStudy (igreen):

Yep, that gives us: 3x + 4 = 5 Now subtract 4 to both sides

OpenStudy (anonymous):

So its B!

OpenStudy (anonymous):

????

OpenStudy (igreen):

Um, no.. Subtract 4 to both sides, what's 5 - 4?

OpenStudy (anonymous):

1! OH I GET IT

OpenStudy (anonymous):

ITS D!

OpenStudy (igreen):

Yep!

OpenStudy (igreen):

3x = 1